Calculate the mean free path of gas molecule in a chamber of 10-7 mm of
mercury pressure. Assuming the molecular diameter to be 3Ȧ. Two grams of
the molecule of the gas occupy 44.8 L at NTP. Consider the temperature of
the chamber to be 273K.
1
Expert's answer
2020-07-21T12:37:57-0400

Pressure "P=133.322\times10^{-7}" Pa

Diameter "d=3\times10^{-10}m"

Temperature is "T=273K"

Now, mean free path "(l)" is

"l=\dfrac{kT}{\sqrt{2}{d^2}P}"

Therefore,

"l=\dfrac{1.38×10^{ 
−23}\times 273}{\sqrt{2}\times 9\times10^{-20}\times 133.332\times 10^{-7}}=2220.31m"